I need to solve this question for my algebra review homework. 5-3(2x-1) is greater than or equal to 6-2x
Answer by stanbon(75887)

5-3(2x-1) >= 6-2x
----
5-6x+3 >= 6-2x
----
8-6x >= 6-2x
----
Add 6x to both sides to get:
2 >= 4x
-----
Reverse te order and divde by 4 to get:
x <= 1/2
==============
Cheers,
Stan H.
=======